# The Effects of Multi-Task Learning on ReLU Neural Network Functions

### v3

@@ 4 Multivariate Multi-Task Neural Network Training
In Section 3, we proved that the univariate functions learned by neural networks trained for multiple tasks can be profoundly different from the functions learned by networks trained for each task separately. In this section, we demonstrate that a similar phenomenon occurs in multivariate settings. Here we analyze neural networks of the form
@@
The above reasoning can be made precise using the concept of exchangeability. A sequence of random variables is exchangeable if their joint distribution remains unchanged for any permutation of the sequence. Intuitively, exchangeability captures the notion that the order of the random variables does not matter. This is naturally consistent with multi-task learning problems, in which the order in which the tasks are enumerated and assigned to the outputs of the network is irrelevant. Therefore, the $T$ tasks can be assigned uniformly at random to the $T$ outputs of the network. This random assignment process induces a distribution on the training data (labels) for each output, denoted by ${\bm{y}}_{\cdot,1},\dots,{\bm{y}}_{\cdot,T}\in\mathbb{R}^{N}$. These vectors are comprised of the $N$ labels for each task corresponding to the inputs ${\bm{x}}_{1},\dots,{\bm{x}}_{N}$. In particular, ${\bm{y}}_{\cdot,1},\dots,{\bm{y}}_{\cdot,T}$ are exchangeable. In the remainder of the discussion, we assume that the vectors ${\bm{y}}_{\cdot,1},\dots,{\bm{y}}_{\cdot,T}$ have been generated according to this procedure.
@@
Let $\mathcal{S}$ be the set of optimal active neurons solving (11), that is, the neurons for which $\|{\bm{v}}^{*}_{k}\|_{2}>0$. For an individual task $s$, consider the objective
@@
where ${\bm{v}}_{k\setminus s}^{*}$ denotes the vector ${\bm{v}}_{k}^{*}$ with its $s^{\text{th}}$ element $v_{ks}^{*}$ excluded, and ${\bm{\Phi}}\in\mathbb{R}^{N\times K}$ is a matrix whose $i,k^{\text{th}}$ entry is ${\bm{\Phi}}_{ik}=\big{(}{\bm{w}}_{k}^{*}\cdot{\bm{x}}_{i}+b_{k}^{*}\big{)}_{+}$. Then with probability at least $1-O(T^{-1/3})$, the regularization term in $H$ is well-defined, and
@@
We provide empirical evidence for the claims presented in this section in Figure 5 on a simple multi-variate dataset. First, we demonstrate the variety of solutions that interpolate this dataset in a single task setting. In contrast, we show that the solutions obtained via multi-task learning with additional random tasks are very similar and often much smoother than those obtained by single-task learning supporting our claim that these solutions are well approximated by a kernel method. We also verify that the optimization (14) is a good approximation for (13). We include additional experiments in Appendix B that demonstrate that these observations hold across multiple trials.
@@
Figure 5: ReLU network interpolation in two-dimensions. The solutions shown were obtained with regularization parameter $\lambda\approx 0$. Top Row – Solutions to single-task training: Figures 5(a), LABEL:, 5(b), LABEL: and 5(c) show solutions to ReLU neural network interpolation (blue surface) of training data (red). The eight data points are located at the vertices of two squares, both centered at the origin. The outer square has side-length two and values of 00 at the vertices. The inner square has side-length one and values of $1$ at the vertices. All three functions interpolate the data and are global minimizers of (2) and (3) when solving for just this task (i.e., $T=1$). Due to the simplicity of this dataset the optimality of the solutions in the first row were confirmed by solving the equivalent convex optimization to (2) developed in Ergen and Pilanci (2021). Bottom Row – Solutions to multi-task training: Figure 5(d) shows the solution to the first output of a multi-task neural network with $T=101$ tasks. The first output is the original task depicted in the first row while the labels for other $100$ tasks are randomly generated i.i.d from a Bernoulli distribution with equal probability for one and zero. Here we show one representative example; more examples are depicted in Appendix B showing that this phenomenon holds across many runs. Figure 5(e) shows the solution to fitting the training data by solving (26) over a fixed set of features learned by the multi-task neural network with $T=100$ random tasks. We observe that unlike the highly variable solutions of single-task optimization problem, the solutions obtained by solving the multi-task optimizations are nearly identical, as one would have for kernel methods. Moreover, the solution obtained by solving (26) is also similar to the solution of the full multi-task training problem with all $T=101$ tasks.
@@ Appendix A Proofs of Main Results
    $v_{1}(x+b_{1})_{+}+v_{2}(-x-b_{1})_{+}=(v_{1}+v_{2})(x+b_{1})_{+}-(v_{1}+v_{2})(x+b_{1}).$
    The term $(v_{1}+v_{2})(x+b_{1})$ can be absorbed into the residual connection, representing the same function with no greater representational cost.
@@
If $N=2$ or $N=3$, then $f_{{\mathcal{D}}}$ is the only solution to (3), so we focus on the case where $N\geq 4$. Suppose that for some $i\in\{2,\dots,N-2\}$, the data points $x_{i-1},x_{i},x_{i+1},x_{i+2}\in\mathbb{R}$ and labels ${\bm{y}}_{i-1},{\bm{y}}_{i},{\bm{y}}_{i+1},{\bm{y}}_{i+2}\in\mathbb{R}^{T}$ satisfy the requirement that
@@
$(x_{i+1}-x_{i}){\bm{y}}_{i-1}-w(x_{i+2}-x_{i+1}){\bm{y}}_{i}+\big{(}(1-w)x_{i}-x_{i+1}+wx_{i+2}\big{)}{\bm{y}}_{i+1}-w(x_{i+1}-x_{i}){\bm{y}}_{i+2}=\bm{0}$ (31)
or equivalently
$\underbrace{[{\bm{y}}_{i-1},{\bm{y}}_{i},{\bm{y}}_{i+1},{\bm{y}}_{i+2}]}_{{\bm{Y}}_{i}\in\mathbb{R}^{T\times 4}}\Bigg{(}\underbrace{\begin{bmatrix}x_{i+1}-x_{i}\\ 0\\ x_{i}-x_{i+1}\\ 0\end{bmatrix}}_{{\bm{a}}_{1}\in\mathbb{R}^{4}}-w\underbrace{\begin{bmatrix}0\\ x_{i+2}-x_{i+1}\\ -x_{i+2}\\ x_{i+1}-x_{i}\end{bmatrix}}_{{\bm{a}}_{2}\in\mathbb{R}^{4}}\Bigg{)}=\bm{0}$ (32)
for some $w>0$. In order for this requirement to be satisfied, it must be the case that ${\bm{Y}}{\bm{a}}_{1}=w{\bm{Y}}{\bm{a}}_{2}$ for some $w>0$, or equivalently, that the matrix ${\bm{U}}={\bm{Y}}[{\bm{a}}_{1},{\bm{a}}_{2}]\in\mathbb{R}^{T\times 2}$ has rank one. Since the rank of any matrix and its Gram matrix are equivalent, this is equivalent to requiring that ${\bm{U}}{\bm{U}}^{\top}\in\mathbb{R}^{T\times T}$ has rank one, or equivalently (because $T>1$), that $\det({\bm{U}}{\bm{U}}^{\top})=0$. Now observe that, based on the definition of the determinant and of the matrix ${\bm{U}}$, the function $\det({\bm{U}}{\bm{U}}^{\top})$ is a real-valued polynomial function of the variables ${\bm{Y}}_{i}=[{\bm{y}}_{i-1},{\bm{y}}_{i},{\bm{y}}_{i+1},{\bm{y}}_{i+2}]\in\mathbb{R}^{T\times 4}$ and $x_{i-1},x_{i},x_{i+1},x_{i+2}\in\mathbb{R}^{4}$. Therefore, as the zero set of a polynomial function (i.e. an algebraic variety), the set of all ${\bm{Y}}_{i}=[{\bm{y}}_{i-1},{\bm{y}}_{i},{\bm{y}}_{i+1},{\bm{y}}_{i+2}]\in\mathbb{R}^{T\times 4}$ and $x_{i-1},x_{i},x_{i+1},x_{i+2}\in\mathbb{R}^{4}$ for which $\det({\bm{U}}{\bm{U}}^{\top})=0$ has Lebesgue measure zero. This is true for any $i=1,\dots,N$, so taking the union over $i=1,\dots,N$, we have that the set of all possible data points $x_{1},\dots,x_{N}\in\mathbb{R}$ and label vectors ${\bm{y}}_{1},\dots,{\bm{y}}_{N}\in\mathbb{R}^{T}$ which admit non-unique solutions has Lebesgue measure zero (as a subset of $\mathbb{R}^{N}\times\mathbb{R}^{T\times N}$). ∎
@@
where ${\bm{y}}=[y_{1},\dots,y_{N}]^{\top}$ and ${\bm{D}}\in\mathbb{R}^{N\times K}$ is a matrix whose $i$, $k^{\textrm{th}}$ entry is ${\bm{D}}_{ik}=\left({\bm{w}}_{k}^{\top}{\bm{x}}_{i}+b_{k}\right)_{+}$, where ${\bm{w}}_{k}$, $b_{k}$ are chosen such that ${\bm{D}}$ is full-rank, which guarantees that the problem above is well-posed. Then $\|{\bm{v}}^{*}\|_{1}\leq\|{\bm{u}}^{*}\|_{1}$. To see this, note that $\|{\bm{u}}^{*}\|_{1}$ is an upper bound on the optimal objective value $S^{*}$ of the neural network interpolation problem (3) for single-task labels $y_{1},\dots,y_{N}$, since (3) allows for optimizing over the ${\bm{w}}_{k},b_{k}$ in addition to the $v_{k}$. Moreover, $S^{*}\geq\|{\bm{v}}^{*}\|_{1}$, since otherwise the optimal parameters for (3) would provide a smaller-than-optimal objective value for (11). Additionally, $\|{\bm{u}}^{*}\|_{1}\leq\|{\bm{D}}^{+}{\bm{y}}\|_{1}$, where ${\bm{D}}^{+}={\bm{D}}^{\top}({\bm{D}}{\bm{D}}^{\top})^{-1}$ is the pseudo-inverse of ${\bm{D}}$, this is because the vector $\tilde{{\bm{u}}}={\bm{D}}^{+}{\bm{y}}$ necessarily satisfies the interpolation constraint in (50). Furthermore,
@@
With these results in hand, take ${\bm{\theta}}^{*}$ to be a set of optimal active neuron parameters as in the statement of Theorem 4.1. For an individual $k$, define $g(v_{ks}):=\left\|\begin{bmatrix}v_{ks}\\ {\bm{v}}_{k\setminus s}^{*}\end{bmatrix}\right\|_{2}=\sqrt{v_{ks}^{2}+\|{\bm{v}}_{k\setminus s}^{*}\|_{2}^{2}}$. Note that $g$ is infinitely differentiable everywhere unless $\|{\bm{v}}_{k\setminus s}^{*}\|_{2}=0$. Since ${\bm{\theta}}^{*}$ is restricted to active neurons, this is only possible if $v_{ks}^{*}\neq 0$, in which case $\frac{(v_{ks}^{*})^{2}}{\|{\bm{v}}_{k\setminus s}^{*}\|_{2}^{2}}=\infty$ (see footnote 6). Therefore, in the event that $\frac{(v_{ks}^{*})^{2}}{\|{\bm{v}}_{k\setminus s}^{*}\|_{2}^{2}}\leq T^{-\beta}$ (for some fixed $10<\beta<$ 1)—which occurs with probability at least $1-(T^{\beta}+1)/T$ by Lemma A.2—we can express $g$ (for this $k$) using Taylor’s theorem as

### v4

@@ 4 Multivariate Multi-Task Neural Network Training
In Section 3, we proved that the univariate-input functions learned by neural networks trained on multiple tasks simultaeously can be profoundly different from the functions learned by networks trained for each task separately. In this section, we demonstrate that a similar phenomenon occurs in multivariate settings. Here we analyze neural networks of the form
@@
The above reasoning can be made precise using the concept of exchangeability. A sequence of random variables is exchangeable if their joint distribution remains unchanged for any finite permutation of the sequence. Intuitively, exchangeability captures the notion that the order of the random variables does not matter. This is naturally consistent with multi-task learning problems, in which the order in which the tasks are enumerated and assigned to the outputs of the network is irrelevant. Therefore, the $T$ tasks can be assigned uniformly at random to the $T$ outputs of the network. This random assignment process induces a distribution on the training data (labels) for each output, denoted by ${\bm{y}}_{\cdot,1},\dots,{\bm{y}}_{\cdot,T}\in\mathbb{R}^{N}$. These vectors are comprised of the $N$ labels for each task corresponding to the inputs ${\bm{x}}_{1},\dots,{\bm{x}}_{N}$. In particular, ${\bm{y}}_{\cdot,1},\dots,{\bm{y}}_{\cdot,T}$ are exchangeable. In the remainder of the discussion, we assume that the vectors ${\bm{y}}_{\cdot,1},\dots,{\bm{y}}_{\cdot,T}$ have been generated according to this procedure.
@@
Let $\mathcal{S}$ be the set of optimal active neurons [^6] A neuron $\eta({\bm{x}})={\bm{v}}({\bm{w}}^{T}{\bm{x}}+b)_{+}$ is active if $\|{\bm{v}}\|_{2}>0$. solving (12), that is, the neurons for which $\|{\bm{v}}^{*}_{k}\|_{2}>0$. For an individual task $s$, consider the objective
@@
where ${\bm{v}}_{k\setminus s}^{*}$ denotes the vector ${\bm{v}}_{k}^{*}$ with its $s^{\text{th}}$ element $v_{ks}^{*}$ excluded, and ${\bm{\Phi}}\in\mathbb{R}^{N\times K}$ is a matrix whose $i,k^{\text{th}}$ entry is ${\bm{\Phi}}_{ik}=\big{(}{\bm{x}}_{i}^{\top}{\bm{w}}_{k}^{*}+b_{k}^{*}\big{)}_{+}$. Then with probability at least $1-O(T^{-1/3})$, the regularization term in $H$ is well-defined, and
@@
We provide empirical evidence for the claims presented in this section in Figure 5 on a simple multi-variate dataset. First, we demonstrate the variety of solutions that interpolate this dataset in a single task setting. In contrast, we show that the solutions obtained via multi-task learning with additional random tasks are very similar and often much smoother than those obtained by single-task learning supporting our claim that these solutions are well approximated by solving a kernel ridge regression problem. We also verify that the optimization (15) is a good approximation for (14). We include additional experiments in Appendix B that demonstrate that these observations hold across multiple trials.
@@
Figure 5: ReLU network interpolation in two-dimensions. The solutions shown were obtained with regularization parameter $\lambda\approx 0$. Top Row – Solutions to single-task training: Figures 5(a), LABEL:, 5(b), LABEL: and 5(c) show solutions to ReLU neural network interpolation (blue surface) of training data (red). The eight data points are located at the vertices of two squares, both centered at the origin. The outer square has side-length two and values of 00 at the vertices. The inner square has side-length one and values of $1$ at the vertices. All three functions interpolate the data and are global minimizers of (2) and (3) when solving for just this task (i.e., $T=1$). Due to the simplicity of this dataset the optimality of the solutions in the first row were confirmed by solving the equivalent convex optimization to (2) developed in Ergen and Pilanci (2021). Bottom Row – Solutions to multi-task training: Figure 5(d) shows the solution to the first output of a multi-task neural network with $T=101$ tasks. The first output is the original task depicted in the first row while the labels for other $100$ tasks are randomly generated i.i.d from a Bernoulli distribution with equal probability for one and zero. Here we show one representative example; more examples are depicted in Appendix B showing that this phenomenon holds across many runs. Figure 5(e) shows the solution to fitting the training data by solving (27) over a fixed set of features learned by the multi-task neural network with $T=100$ random tasks. We observe that unlike the highly variable solutions of single-task optimization problem, the solutions obtained by solving the multi-task optimizations are nearly identical, as one would have for kernel methods. Moreover, the solution obtained by solving (27) is also similar to the solution of the full multi-task training problem with all $T=101$ tasks. Refer to caption
@@ Appendix A Proofs of Main Results
    $v_{1}(x+b_{1})_{+}+v_{2}(-x-b_{1})_{+}=(v_{1}+v_{2})(x+b_{1})_{+}-v_{2}(x+b_{1}).$
    The term $-v_{2}(x+b_{1})$ can be absorbed into the residual connection, representing the same function with no greater representational cost.
@@
If $N=2$ or $N=3$, then $f_{{\mathcal{D}}}$ is the only solution to (3), so we focus on the case where $N\geq 4$. Suppose that for some $i=2,\dots,N-2$, the data points $x_{i-1},x_{i},x_{i+1},x_{i+2}\in\mathbb{R}$ and labels ${\bm{y}}_{i-1},{\bm{y}}_{i},{\bm{y}}_{i+1},{\bm{y}}_{i+2}\in\mathbb{R}^{T}$ satisfy the requirement that
@@
$\underbrace{[{\bm{y}}_{i-1},{\bm{y}}_{i},{\bm{y}}_{i+1},{\bm{y}}_{i+2}]}_{{\bm{Y}}_{i}\in\mathbb{R}^{T\times 4}}\Bigg{(}\underbrace{\begin{bmatrix}\frac{1}{x_{i}-x_{i-1}}\\ \frac{1}{x_{i+1}-x_{i}}-\frac{1}{x_{i}-x_{i-1}}\\ \frac{1}{x_{i+1}-x_{i}}\\ 0\end{bmatrix}}_{{\bm{a}}_{1}\in\mathbb{R}^{4}}-w\underbrace{\begin{bmatrix}0\\ \frac{1}{x_{i+1}-x_{i}}\\ -\frac{1}{x_{i+2}-x_{i+1}}-\frac{1}{x_{i+1}-x_{i}}\\ \frac{1}{x_{i+2}-x_{i+1}}\end{bmatrix}}_{{\bm{a}}_{2}\in\mathbb{R}^{4}}\Bigg{)}=\bm{0}$ (32)
or equivalently, that ${\bm{Y}}{\bm{a}}_{1}=w{\bm{Y}}{\bm{a}}_{2}$ for some $w>0$. If this condition is satisfied, the matrix ${\bm{U}}={\bm{Y}}[{\bm{a}}_{1},{\bm{a}}_{2}]\in\mathbb{R}^{T\times 2}$ has rank at most one, as does its Gram matrix ${\bm{U}}{\bm{U}}^{\top}$, which implies (because $T>1$) that $\det({\bm{U}}{\bm{U}}^{\top})=0$. Observe that $\det({\bm{U}}{\bm{U}}^{\top})$ is a real-valued non-trivial rational function of the variables $x_{i-1},x_{i},x_{i+1},x_{i+2}\in\mathbb{R}$ and the entries of ${\bm{Y}}_{i}=[{\bm{y}}_{i-1},{\bm{y}}_{i},{\bm{y}}_{i+1},{\bm{y}}_{i+2}]\in\mathbb{R}^{T\times 4}$. The zero set of this rational function is contained in the zero set of its polynomial numerator (an algebraic variety), and thus has Lebesgue measure zero in $\mathbb{R}^{4}\times\mathbb{R}^{T\times 4}$, which implies that the set of possible $x_{i-1},x_{i},x_{i+1},x_{i+1},{\bm{Y}}_{i}$ which satisfy the original condition for any $w>0$ does as well. This holds for any $i=2,\dots,N-2$, so the union of all such sets over $i=2,\dots,N-2$ also has Lebesgue measure zero in $\mathbb{R}^{N}\times\mathbb{R}^{T\times N}$. ∎
@@
where ${\bm{y}}=[y_{1},\dots,y_{N}]^{\top}$ and ${\bm{D}}\in\mathbb{R}^{N\times K}$ is a matrix whose $i$, $k^{\textrm{th}}$ entry is ${\bm{D}}_{ik}=\left({\bm{w}}_{k}^{\top}{\bm{x}}_{i}+b_{k}\right)_{+}$, where ${\bm{w}}_{k}$, $b_{k}$ are chosen such that ${\bm{y}}\in\text{range}({\bm{D}})$, which guarantees that the problem above is well-posed. Then $\|{\bm{v}}^{*}\|_{1}\leq\|{\bm{u}}^{*}\|_{1}$. To see this, note that $\|{\bm{u}}^{*}\|_{1}$ is an upper bound on the optimal objective value $S^{*}$ of the neural network interpolation problem (3) for single-task labels $y_{1},\dots,y_{N}$, since (3) allows for optimizing over the ${\bm{w}}_{k},b_{k}$ in addition to the $v_{k}$. Moreover, $S^{*}\geq\|{\bm{v}}^{*}\|_{1}$, since otherwise the optimal parameters for (3) would provide a smaller-than-optimal objective value for (12). Additionally, $\|{\bm{u}}^{*}\|_{1}\leq\|{\bm{D}}^{+}{\bm{y}}\|_{1}$, where ${\bm{D}}^{+}={\bm{D}}^{\top}({\bm{D}}{\bm{D}}^{\top})^{-1}$ is the pseudo-inverse of ${\bm{D}}$, this is because the vector $\tilde{{\bm{u}}}={\bm{D}}^{+}{\bm{y}}$ necessarily satisfies the interpolation constraint in (50). Furthermore,
@@
With these results in hand, take ${\bm{\theta}}^{*}$ to be a set of optimal active neuron parameters as in the statement of Theorem 4.1. For an individual $k$, define $g(v_{ks}):=\left\|\begin{bmatrix}v_{ks}\\ {\bm{v}}_{k\setminus s}^{*}\end{bmatrix}\right\|_{2}=\sqrt{v_{ks}^{2}+\|{\bm{v}}_{k\setminus s}^{*}\|_{2}^{2}}$. Note that $g$ is infinitely differentiable everywhere unless $\|{\bm{v}}_{k\setminus s}^{*}\|_{2}=0$. Since ${\bm{\theta}}^{*}$ is restricted to active neurons, this is only possible if $v_{ks}^{*}\neq 0$, in which case $\frac{(v_{ks}^{*})^{2}}{\|{\bm{v}}_{k\setminus s}^{*}\|_{2}^{2}}=\infty$ (see footnote 7). Therefore, in the event that $\frac{(v_{ks}^{*})^{2}}{\|{\bm{v}}_{k\setminus s}^{*}\|_{2}^{2}}\leq T^{-\beta}$ (for some fixed $10<\beta<$ 1)—which occurs with probability at least $1-(T^{\beta}+1)/T$ by Lemma A.2—we can express $g$ (for this $k$) using Taylor’s theorem as